OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (mr.math):

\[x^2+2x-35=(x+7)(x-5)\]

OpenStudy (anonymous):

what factoring techniques do you know?

OpenStudy (lilg132):

use formula

OpenStudy (mr.math):

If \((x+7)(x-5)=0\), then \(x+7=0 \implies x=-7\) or \(x-5=0 \implies x=5\).

OpenStudy (anonymous):

thanks